One year ago, Allison purchased 350 shares of a mutual fund which has a front-end load of 5.25 percent. The NAV at the time of purchase was $30. Today, the NAV is $33. The fund had total annual expenses of 1.65 percent. There were no fund distributions this past year. What is Allison's rate of return for the year?

Group of answer choices

4.76 percent

-1.10 percent

-1.04 percent

4.23 percent

5.00 percent

You want to buy 1,200 shares of a mutual fund that has an NAV of $27.80. The fund charges a 3.75 percent front-end load. How much will you have to spend to make this purchase?

$34,659.74

$35,913.00

$38,445.60

$37,333.33

$36,953.37
Answer #1

1). Purchase Price = NAV at time of purchase / (1 - front-end load)

= $30 / (1 - 0.0525) = $30 / 0.9475 = $31.66

Rate of Return = [Current NAV / Purchase Price] - 1

= [$33 / $31.66] - 1 = 1.04225 - 1 = 0.04225, or 4.23%

So, 4th option is correct.

2). Purchase Price = NAV at time of purchase / (1 - front-end load)

= $27.80 / (1 - 0.0375) = $27.80 / 0.9625 = $28.88

Amount of Investment = Purchase Price * No. of shares outstanding

= $28.88 * 1,200 = $34,659.74

So, 1st option is correct.